OpenStudy (anonymous):

Is the equation (true), (false), or (open)? Explain. 9p + 8 = 10p + 7

OpenStudy (anonymous):

p=1

OpenStudy (anonymous):

9p + 8 = 10p + 7 subtract 9p from both sides 9p - 9p + 8 = 10p - 9p + 7 8 = p + 7 subtract 7 from both sides 8 - 7 = p + 7 -7 1 = p it is true because on substituting p= 1 we find LHS = RHS